Find the sum of the first thirty terms of the following arithmetic series:
16+25+34+43...

Found 2 solutions by stanbon, Fombitz:
Answer by stanbon(57219)

Find the sum of the first thirty terms of the following arithmetic series:
16+25+34+43...
-------------
Find the 30th term:
a(30) = a(1)+29(9)
a(30) = 16+ 261 = 277
-------------------------
S(30) = (30/2)(16+277) = 4395
================================
Cheers,
Stan H.

Answer by Fombitz(13828)

Your series looks like
1st term : 16=16=16+9(1-1)
2nd term : 25=16+9=16+9(2-1)
3rd term : 34=16+9+9=16+9(3-1)
4th term : 43=16+9+9+9=16+9(4-1)
.
.
.
30th term : 261=16+9(29)=16+9(30-1)
The sum of all those terms would be then the sum of 30*16 since there are 30 sixteens added plus 9 times the sum from 1 to 29.
The sum from 1 to n equals n(n+1)/2.
The sum from 1 to 29 equals 29(30)/2=435.
The total sum is then 30*16 (480) plus 9*435 (3915) and equals 4395.